OpenStudy (anonymous):

Find the derivative: f(x)=ln(sqrt((x^2+1)/x(2x^3-1)^2))

OpenStudy (anonymous):

\[\ln\frac{ \sqrt{x^2+1} }{ x(2x^3-1)^2}\]

OpenStudy (anonymous):

derivative of ln(u)= u'/u. so use the quotient rule to take the derivative and divide it by your equation excluding the ln. I would do it but that is one nasty derivative. once you get an answer, you can check it in the web, wolframalpha